Solve the following equations: 2x+9(root over)=13-x
Answer by mananth(16949)

sqrt2x+9 = 13-x
square the equation
2x+9 = (13-x)^2
2x+9=169-26x+x^2
x^2-28x+160=0
x^2-20x-8x+160
x(x-20)-8(x-20)
(x-20)(x-8)
x= 20 OR 8